Helping my daughter on Algebra. It's been 20 years.
1/2 x = 3 5/6
what is x, and how did you do it?
also, 2/3 x = 24
don't understand equation to get to x.
please help.

Answer by stanbon(75887)

1/2 x = 3 5/6
Change the mixed number to a fraction:
(1/2)x = 23/6
---
Multiply both sides of the equality by 2 to get:
x = 2(23/6)
Cancel the common factor of 2 in the numerator and the denominator to get:
x = 23/3
Change to a mixed number
x = 7 2/3
=============================================================
2/3 x = 24
Multiply both sides by 3/2 to get:
x = (3/2)24
Cancel the common factor of 2 to get:
x = 3*12
x = 36
===============
Cheers,
Stan H.